# GENIE Collaboration

### v1

@@ Abstract
We present the first global tuning, using genie, of four transverse kinematic imbalance measurements of neutrino-hydrocarbon scattering, both with and without pion final states, from the T2K and MINERvA experiments. As a proof of concept, we have simultaneously tuned the initial state and final-state interaction models (SF-LFG and hA, respectively), producing a new effective theory that more accurately describes the data.
preprint: APS/123-QED
I Introduction
Neutrinos play a central role in advancing our understanding of physics and addressing fundamental inquiries in contemporary science. The Hyper-Kamiokande experiment [1] aims to conduct precise measurements of charge parity violation within the neutrino sector, a phenomenon believed to be closely linked to the observed matter-antimatter asymmetry in the universe. Similarly, the Deep Underground Neutrino Experiment (DUNE) [2, 3, 4, 5] promises to elucidate the neutrino mass ordering. Owing to their substantial scale, these forthcoming experiments are ideally suited for exploring novel physics phenomena, such as neutrino-less double beta decay. Whether seeking to ascertain Standard Model parameters or probing for exotic phenomena, significant enhancements in both theoretical frameworks and computational simulations of neutrino-nucleus interactions are imperative. As neutrinos interact with nucleons inside the nucleus, the interaction is subject to nuclear effects, namely the nucleon initial state and final-state interactions (FSIs). These effects are not fully understood yet and can alter the final event topology by emission or absorption of pions. Notably, the neutrino flux at DUNE yields a comparable proportion of events with and without pions in the final state, highlighting the pressing need for a generator capable of accurately describing both event topologies.
@@ I Introduction
This paper is structured as follows: In Sec. II, we review the TKI measurements and genie models. In Sec. III, we detail the tuning considerations and procedures. Results are summarized in Sec. IV, highlighting how the data-MC discrepancy in the MINERvA-$\pi^{0}$ TKI measurement [12] is resolved while maintaining good data-MC agreement elsewhere. We conclude in Sec. V.
II TKI measurements and genie models
@@
TKI is a methodology focused on the conservation of momentum in neutrino interactions. In essence, it involves examining the imbalance between the observed transverse momentum of the final-state particles and the expected transverse momentum from neutrino interactions with free nucleons [10, 11]. This “kinematic mismatch,” together with its longitudinal and three-dimensional variations [13, 14], and the resulting asymmetry [15], has been a crucial set of observables, establishing a pathway to extract valuable information about the participating particles and the underlying nuclear processes. Recent experimental results from neutrino experiments such as T2K [16, 17], MINERvA [18, 19, 12], and MicroBooNE [20, 21, 22, 23, 24], as well as electron scattering experiments such as CLAS [25], highlight the efficacy of TKI.
@@
In neutrino scattering off a free nucleon, the sum of the transverse components of the final products is expected to be zero, visualized through a back-to-back configuration between the final-state lepton and hadronic system in the plane transverse to the neutrino direction. Hence, in a neutrino interaction with a nucleus, the transverse momentum imbalance, $\delta p_{\textrm{T}}$ [11], results from intranuclear dynamics, including Fermi motion and FSIs. The deviation from being back-to-back is quantified by the coplanarity angle $\delta\phi_{\textrm{T}}$, and the transverse boosting angle, $\delta\alpha_{\textrm{T}}$ [11], indicates the direction of $\delta p_{\textrm{T}}$ within the transverse plane. Furthermore, analyzing the energy and longitudinal momentum budget [13, 14] enables the conversion of $\delta p_{\textrm{T}}$ to the emulated nucleon momentum, $p_{\textrm{N}}$, providing insight into the Fermi motion; this conversion amounts to a correction on the order of $\mathcal{O}(20\%)$ [26]. In the absence of FSIs, $\delta\alpha_{\textrm{T}}$ remains uniform (except for second-order effects, such as variations in the center-of-mass energy), given the isotropic nature of the initial nucleon motion. However, as the final products propagate through the nuclear medium, they experience FSIs, thereby disturbing the isotropy and the Fermi motion peak of the $\delta\alpha_{\textrm{T}}$ and $p_{\textrm{N}}$ ($\delta p_{\textrm{T}}$) distributions, respectively. Hence, $\delta p_{\textrm{T}}$ and $p_{\textrm{N}}$ elucidate the Fermi motion details, while $\delta\alpha_{\textrm{T}}$ characterizes the FSI strength—crucial for elucidating medium effects in neutrino interactions. Moreover, a notable advantage of these observables is their minimal dependence on neutrino energy [11]. The double TKI variable, $\delta p_{\textrm{TT}}$ [10], is the projection of $\delta\vec{p}_{\textrm{T}}$ along the axis perpendicular to the lepton scattering plane (hence “double”). In addition to its use for extracting neutrino-hydrogen interactions [10, 27], it has also been applied to study nuclear effects in neutrino pion productions [12, 17]. Its equivalent in pionless production, $\delta p_{\textrm{Tx}}$, has been proposed and studied together with its orthogonal companion, $\delta p_{\textrm{Ty}}$, in MINERvA [19].
@@
This work surveyed four TKI data sets: T2K $0\pi$ [16], T2K $\pi^{+}$ [17], MINERvA $0\pi$ [18, 19], and MINERvA $\pi^{0}$ [12] measurements. All four measurements require the presence of one CC muon and at least one proton in the final state. While the T2K-$\pi^{+}$ measurement requires exactly one $\pi^{+}$, and MINERvA-$\pi^{0}$ requires at least one $\pi^{0}$, the other two datasets require the absence of any pions. The definitions of the phase spaces for these signals are shown in Table 1.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 1: $\delta\alpha_{\textrm{T}}$ measurements decomposed in interaction types, compared to the nominal tune, G24-0.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 2: Similar to Fig. 2 but for the $p_{\textrm{N}}$ ($\delta p_{\textrm{T}}$ for T2K-$0\pi$) measurements.
Variables Cut Values ($p$ in $\textrm{GeV}/c$)
T2K-$0\pi$ [16]
$\vec{p}_{\mu}$ cut $0.25<p_{\mu}$, $\cos\theta_{\mu}>-0.6$
$\vec{p}_{\textrm{p}}$ cut $1.00.45<p_{\text{p}}<$ 1.0 , $\cos\theta_{\text{p}}>0.4$
T2K-$\pi^{+}$ [17]
$\vec{p}_{\mu}$ cut $70.25<p_{\mu}<$ 7 , $\theta_{\mu}<70^{\circ}$
$\vec{p}_{\textrm{p}}$ cut $1.20.45<p_{\text{p}}<$ 1.2 , $\theta_{\text{p}}<70^{\circ}$
$\vec{p}_{\pi}$ cut $1.20.15<p_{\pi}<$ 1.2, $\theta_{\pi}<70^{\circ}$
MINERvA-$0\pi$ [18, 19]
$\vec{p}_{\mu}$ cut $101.5<p_{\mu}<$ 10 , $\theta_{\mu}<20^{\circ}$
$\vec{p}_{\textrm{p}}$ cut $1.20.45<p_{\text{p}}<$ 1.2 , $\theta_{\text{p}}<70^{\circ}$
MINERvA-$\pi^{0}$ [12]
$\vec{p}_{\mu}$ cut $201.5<p_{\mu}<$ 20 , $\theta_{\mu}<25^{\circ}$
$\vec{p}_{\textrm{p}}$ cut $0.45<p_{\text{p}}$
Table 1: Phase space cuts for the signals of the TKI measurements.
@@
The genie models and tunes are detailed in the Comprehensive Model Configurations. Both the aforementioned G24-0 and the widely-used G18_10a_02_11b tunes derive from the same tuning process [7, 8], yet G24-0 distinguishes itself by three aspects: 1) incorporating a new initial state, the correlated and spectral-function-like local Fermi gas (SF-LFG) [35, 36, 37]; 2) using $z$-expansion axial-vector form factor [38] rather than the dipole form factor of the Valencia model in QE processes [39] and 3) replacing the Valencia model [40] in 2p2h processes with the SuSAv2 Model [41]. The other model components are given in Table  2. Given its simplicity and interpretability, the INTRANUKE hA FSI model (hA for short) emerges as a strong candidate for an effective model. Since the TKI measurements are sensitive to the initial state and FSI, this work investigates a partial tune of these two processes based on G24-0, in an effort to derive an effective model describing neutrino-nuclear pion production, while maintaining the efficacy with the pionless production.
Simulation domain Model
Nuclear State SF-LFG [35, 36, 37]
QE Valencia [39]
2p2h SuSAv2 [41]
QE $\Delta S=1$ Pais [42]
QE $\Delta C=1$ Kovalenko [43]
Resonance (RES) Berger-Sehgal [44]
Shallow/Deep inelastic scattering (SIS/DIS) Bodek-Yang [45]
DIS $\Delta C=1$ Aivazis-Tung-Olness [46]
Coherent $\pi$ production Berger-Sehgal [47]
Hadronization AGKY [48]
FSI INTRANUKE hA [49]
@@
Table 2: Model components of the nominal G24-0 tune. Processes with non-trivial $\Delta S$ and $\Delta C$ are those with strangeness and charm production, respectively.
III Tuning Initial-State and FSI models on TKI Data
@@
This study adopts the tuning procedure outlined in Ref. [9], utilizing $N_{\textrm{par}}$ input model parameters. The objective is to identify the extremal point—that is, the optimal set of parameter values—within the parameter space that minimizes the $\chi^{2}$ between model predictions and data. Given the high dimensionality of the parameter space, conducting a brute-force, point-by-point scan on a grid is infeasible. Therefore, a sufficient number of points is randomly sampled. Each point is used to run a full simulation. The simulation output for each data observable bin is then parameterized by Professor [50] with a polynomial of the model parameters of a chosen order, $N_{\textrm{ord}}$. The minimal number of parameter sets ($N_{\textrm{s}}$) required for tuning $N_{\textrm{par}}$ parameters with an order $N_{\textrm{ord}}$ polynomial is determined by the combination formula,
@@
Given that $N_{\textrm{s}}$ increases factorially, caution is essential when augmenting the number of parameters or the polynomial degree. Given that our tuning incorporates all $2$ SF-LFG and $12$ hA parameters parameters (see Sec. III.2), a degree-$4$ parameterization necessitates over 6,000 generations. Hence, only up to order $4$ polynomials are explored, and it has shown to reproduce the MC predictions well as shown in Fig. 3.
Figure 3: Fractional difference for the bin-by-bin cross sections between MC truth and the parameterized approximation with order-4 polynomials, both in the Norm-Shape (NS) [51, 52] space with AllPar. See later text for the definition of AllPar. The residual has a mean of $0.003$ and a standard deviation of $0.016$.
Furthermore, to circumvent Peelle’s Pertinent Puzzle [53], the Norm-Shape (NS) transformation prescription [51, 52] is applied. The extremal point is then found from a minimization of $\chi^{2}$ between this NS-transformed polynomial approximation and NS-transformed data. In the minimization, priors, usually based on systematic uncertainty, can be imposed on each parameter to penalize it from getting too far from its default value. The following subsections elaborate on the specific choice of measurement observables and model parameters in this work.
III.1 Data Observables
The observables for the reported differential cross sections across the four TKI measurements—T2K-$0\pi$, T2K-$\pi^{+}$, MINERvA-$0\pi$, and MINERvA-$\pi^{0}$—are detailed in Table 3. To identify the most sensitive variables, various combinations were evaluated for tuning purposes. A total of 26 combinations were assessed, with the superset comprising $\delta\alpha_{\textrm{T}}$, $\delta p_{\textrm{T}}$, $\delta\phi_{\textrm{T}}$, $p_{\textrm{N}}$, and $\delta p_{\textrm{TT}}$ (see Table 6 in Appendix .1 for all combinations). Each combination was evaluated such that non-selected variables—including proton momentum and angle ($p_{\text{p}}$ and $\theta_{\text{p}}$[18], as well as $\delta p_{\textrm{Tx}}$ and $\delta p_{\textrm{Ty}}$ [19] in MINERvA-$0\pi$—were used solely for validation; the model is expected to accurately predict these variables post-tuning. When constructing the combinations, it was observed that $\delta\phi_{\textrm{T}}$ is strongly dependent on neutrino beam energy [11], which suppresses its sensitivity to nuclear effects. Additionally, correlations exist between various observables, notably between $\delta p_{\textrm{T}}$ and $p_{\textrm{N}}$. Therefore, determining the optimal combination is not straightforward; this study employs a criterion based on the $\chi^{2}$ with the complete (tuned plus validation) observable set (Fig. 4).

### v2

# Oxford OX1 3RH, UK

@@ Abstract
We present the first combined tuning, using genie, of four transverse kinematic imbalance measurements of neutrino-hydrocarbon scattering, both with and without pion final states, from the T2K and MINERvA experiments. As a proof of concept, we have simultaneously tuned the initial state and final-state interaction models (SF-CFG and hA, respectively), producing a new effective model that more accurately describes the data.
arxiv: 2404.08510
1 Introduction
Neutrinos play a central role in advancing our understanding of physics and addressing fundamental inquiries in contemporary science. The Hyper-Kamiokande experiment Hyper-Kamiokande:2018ofw aims to conduct precise measurements of charge parity violation within the neutrino sector, a phenomenon believed to be closely linked to the observed matter-antimatter asymmetry in the universe. Similarly, the Deep Underground Neutrino Experiment (DUNE) DUNE:2016hlj ; DUNE:2015lol ; DUNE:2016evb ; DUNE:2016rla ; DUNE:2021tad promises the same, in addition to elucidating the neutrino mass ordering. Whether seeking to ascertain Standard Model parameters or probing for exotic phenomena, significant enhancements in both theoretical frameworks and computational simulations of neutrino-nucleus interactions are imperative. As neutrinos interact mainly with nucleons inside the nucleus, the interaction is subject to nuclear effects, namely the nucleon initial state and final-state interactions (FSIs). These effects are difficult to calculate and can alter the final event topology by changing the number of final state pions with respect to the interactions on free nucleons. Notably, the neutrino flux at DUNE yields a comparable proportion of events with and without pions in the final state, highlighting the pressing need for a generator capable of accurately describing both event topologies.
@@ 1 Introduction
This paper is structured as follows: In Sec. 2 and Sec. 3, we review the TKI measurements and genie models respectively. In Sec. 4, we detail the tuning considerations and procedures. Results are summarized in Sec. 5, highlighting how the data-MC discrepancy in the MINERvA-$\pi^{0}$ TKI measurement MINERvA:2020anu is resolved while maintaining good data-MC agreement elsewhere. We conclude in Sec. 6.
2 TKI measurements
@@
TKI is a methodology based on the conservation of momentum in neutrino interactions. In essence, it involves quantifying the imbalance between the observed transverse momentum of the final-state particles and the expected transverse momentum from neutrino interactions with free nucleons Lu:2015hea ; Lu:2015tcr . This “kinematic mismatch” together with its longitudinal and three-dimensional variations Furmanski:2016wqo ; Lu:2019nmf , and the derived asymmetry Cai:2019jzk , has been a crucial set of observables, establishing a pathway to extract valuable information about the participating particles and the underlying nuclear processes. Recent experimental results from neutrino experiments such as T2K T2K:2018rnz ; T2K:2021naz , MINERvA MINERvA:2018hba ; MINERvA:2019ope ; MINERvA:2020anu ; MINERvA:2021csy , and MicroBooNE MicroBooNE:2022emb ; MicroBooNE:2023cmw ; MicroBooNE:2023tzj ; MicroBooNE:2023wzy ; MicroBooNE:2024tmp , as well as electron scattering experiments such as CLAS CLAS:2021neh , highlight the efficacy of TKI.
Figure 1: Schematic illustration of the TKI variables. Diagram taken from Ref. Lu:2015tcr .
@@
In neutrino scattering off a free nucleon, the sum of the transverse components of the final products is expected to be zero, visualized through a back-to-back configuration between the final-state lepton and hadronic system in the plane transverse to the neutrino direction. Hence, in a neutrino interaction with a nucleus, the transverse momentum imbalance, $\delta p_{\textrm{T}}$ Lu:2015tcr , results from intranuclear dynamics, including Fermi motion and FSIs as shown in Fig. 1. The deviation from being back-to-back is quantified by the coplanarity angle $\delta\phi_{\textrm{T}}$ Lu:2015tcr , while the transverse boosting angle, $\delta\alpha_{\textrm{T}}$ Lu:2015tcr , represents the direction of $\delta p_{\textrm{T}}$ within the transverse plane. Furthermore, analyzing the energy and longitudinal momentum budget Furmanski:2016wqo ; Lu:2019nmf enables the conversion of $\delta p_{\textrm{T}}$ to the emulated nucleon momentum, $p_{\textrm{N}}$, providing further insight into the Fermi motion; this conversion amounts to a correction on the order of $\mathcal{O}(20\%)$ Yang:2023dxk . With one-body currents in the absence of FSIs, $\delta\alpha_{\textrm{T}}$ remains uniform (except for second-order effects, such as variations in the center-of-mass energy), given the isotropic nature of the initial nucleon motion. However, as the final products propagate through the nuclear medium, they experience FSIs, thereby disturbing the isotropy and the Fermi motion peak of the $\delta\alpha_{\textrm{T}}$ and $p_{\textrm{N}}$ ($\delta p_{\textrm{T}}$) distributions, respectively. Hence, $\delta p_{\textrm{T}}$ and $p_{\textrm{N}}$ elucidate the Fermi motion details, while $\delta\alpha_{\textrm{T}}$ characterizes the FSI strength—crucial for understanding medium effects in neutrino interactions. A notable advantage of these observables is their minimal dependence on neutrino energy Lu:2015tcr . Moreover, the double TKI variable, $\delta p_{\textrm{TT}}$ Lu:2015hea , is the projection of $\delta\vec{p}_{\textrm{T}}$ along the axis perpendicular to the lepton scattering plane (hence “double”). In addition to its use for extracting neutrino-hydrogen interactions Lu:2015hea ; Hamacher-Baumann:2020ogq , it has also been applied to study nuclear effects in neutrino pion productions MINERvA:2020anu ; T2K:2021naz . Its equivalent in pionless production, $\delta p_{\textrm{Tx}}$, has been proposed and studied together with its orthogonal companion, $\delta p_{\textrm{Ty}}$, in MINERvA MINERvA:2019ope .
@@
This work surveyed four TKI data sets: T2K $0\pi$ T2K:2018rnz , T2K $\pi^{+}$ T2K:2021naz , MINERvA $0\pi$ MINERvA:2018hba ; MINERvA:2019ope , and MINERvA $\pi^{0}$ MINERvA:2020anu measurements. All four measurements require the presence of one CC muon and at least one proton in the final state. While the T2K-$\pi^{+}$ measurement requires exactly one $\pi^{+}$, and MINERvA-$\pi^{0}$ requires at least one $\pi^{0}$, the other two datasets require the absence of any pions. The definitions of the kinematic cuts for these samples are summarised in Table 1.
Variables Cuts ($p$ in $\textrm{GeV}/c$)
T2K-$0\pi$ T2K:2018rnz
$\vec{p}_{\mu}$ $0.25<p_{\mu}$, $\cos\theta_{\mu}>-0.6$
$\vec{p}_{\textrm{p}}$ $1.00.45<p_{\text{p}}<$ 1.0 , $\cos\theta_{\text{p}}>0.4$
T2K-$\pi^{+}$ T2K:2021naz
$\vec{p}_{\mu}$ $70.25<p_{\mu}<$ 7 , $\theta_{\mu}<70^{\circ}$
$\vec{p}_{\textrm{p}}$ $1.20.45<p_{\text{p}}<$ 1.2 , $\theta_{\text{p}}<70^{\circ}$
$\vec{p}_{\pi}$ $1.20.15<p_{\pi}<$ 1.2, $\theta_{\pi}<70^{\circ}$
MINERvA-$0\pi$ MINERvA:2018hba ; MINERvA:2019ope
$\vec{p}_{\mu}$ $101.5<p_{\mu}<$ 10 , $\theta_{\mu}<20^{\circ}$
$\vec{p}_{\textrm{p}}$ $1.20.45<p_{\text{p}}<$ 1.2 , $\theta_{\text{p}}<70^{\circ}$
MINERvA-$\pi^{0}$ MINERvA:2020anu
$\vec{p}_{\mu}$ $201.5<p_{\mu}<$ 20 , $\theta_{\mu}<25^{\circ}$
$\vec{p}_{\textrm{p}}$ $0.45<p_{\text{p}}$
Table 1: Kinematic cuts for the samples of the TKI measurements.
@@
3 genie model selection
genie uses formatted strings to uniquely define configurations, which are usually called tunes. Regardless of the name, they can refer to configurations either before or after an actual tuning procedure.
Free nucleon tuning effort in Ref. GENIE:2021zuu has produced a tune, G18_10a_02_11b, with better description of bubble chamber data. This is the starting point of the model we want to tune. The first effect to apply to this model is adding the proper initial state. Among the available genie initial states, the better choice is the spectral-function-like Correlated Fermi gas (SF-CFG) sfcfg-talk ; sfcfg-GitHubCommit ; GENIE:2021npt because it has the most advanced physics and also has the largest number of degrees of freedom that can be adjusted by the tuning procedure. More details on the SF-CFG can be found in Ref. GENIE:2021npt .
Further improvements to the configuration to be tuned include: 1) using $z$-expansion axial-vector form factor Hill:2010yb rather than the dipole form factor of the Valencia model in QE processes Nieves:2004wx ; and 2) replacing the Valencia model Nieves:2011pp in 2p2h processes with the SuSAv2 Model Gonzalez-Jimenez:2014eqa since it covers larger region of the $q^{0}$, $q^{3}$ phase space.
Given its simplicity and interpretability, the INTRANUKE hA FSI model (hA for short) emerges as the FSI candidate for the model to be tuned. The complete list of the model components are given in Table  2. In the genie code, this configuration is identified as G24_20i_00_000 and it is currently the default genie configuration for a number of LAr based experiments. As G24_20i_00_000 is widely used in the paper, we will call it G24-0 for simplicity.
Simulation component Model
Nuclear state SF-CFG sfcfg-talk ; sfcfg-GitHubCommit ; GENIE:2021npt
QE Valencia Nieves:2004wx
2p2h SuSAv2 Gonzalez-Jimenez:2014eqa
QE $\Delta S=1$ Pais Pais:1971er
QE $\Delta C=1$ Kovalenko Kovalenko:1990zi
Resonance (RES) Berger-Sehgal Berger:2007rq
Shallow/Deep inelastic scattering (SIS/DIS) Bodek-Yang Bodek:2002vp
DIS $\Delta C=1$ Aivazis-Tung-Olness Aivazis:1991fy
Coherent $\pi$ production Berger-Sehgal Berger:2008xs
Hadronization AGKY Yang:2009zx
FSI INTRANUKE hA Andreopoulos:2015wxa
@@
Table 2: Model components of G24-0. Processes with non-trivial $\Delta S$ and $\Delta C$ are those with strangeness and charm production, respectively.
Comparing the genie predictions for G24-0 against TKI datasets, Figs. 3 and 3, shows that the model fails to accurately describe the MINERvA $\pi^{0}$ MINERvA:2020anu TKI measurement.
Figure 2: $\delta\alpha_{\textrm{T}}$ measurements decomposed in interaction types, compared to G24-0 prediction.
Figure 3: Similar to Fig. 3 but for the $p_{\textrm{N}}$ (T2K-$\pi^{+}$, MINERvA-$0\pi$ and MINERvA-$\pi^{0}$) and $\delta p_{\textrm{T}}$ (T2K-$0\pi$) measurements.
4 Tuning initial-state and FSI models on TKI data
@@
This study adopts the tuning procedure outlined in Ref. GENIE:2022qrc , utilizing $N_{\textrm{par}}$ model parameters. The objective is to identify the best fit—that is, the optimal set of parameter values—within the parameter space that minimizes the $\chi^{2}$ between model predictions and data. Given the high dimensionality of the parameter space, conducting a brute-force, point-by-point scan on a grid is infeasible. Therefore, a sufficient number of points is randomly sampled. Each point is used to run a full simulation. The simulation output for each data observable bin is then parameterized using Professor Buckley:2009bj with a polynomial of the model parameters of a chosen order, $N_{\textrm{ord}}$. The minimal number of points ($N_{\textrm{s}}$) required for tuning $N_{\textrm{par}}$ parameters with an order $N_{\textrm{ord}}$ polynomial is determined by the combination formula,
@@
Given that $N_{\textrm{s}}$ increases factorially, caution is essential when augmenting the number of parameters or the polynomial degree. Given that our tuning incorporates all $2$ SF-CFG and $12$ hA parameters (see Sec. 4.2), a degree-$4$ parameterization necessitates over 6,000 generations. Hence, only up to order $4$ polynomials are explored, and it has shown to reproduce the MC predictions well, as shown in Fig. 4.
Figure 4: Fractional difference for the bin-by-bin cross sections between MC truth and the parameterized approximation with order-4 polynomials, both in the Norm-Shape (NS) DAgostini:1993arp ; Hanson:2005mrg space with AllPar. See Table 4 for the definition of AllPar. The residual has a mean of $0.003$ and a standard deviation of $0.016$.
Furthermore, to circumvent Peelle’s Pertinent Puzzle PPP_FNL , the Norm-Shape (NS) transformation prescription DAgostini:1993arp ; Hanson:2005mrg is applied. The extremal point is then found from a minimization of $\chi^{2}$ between this NS-transformed polynomial approximation and NS-transformed data. In minimization, priors, usually based on systematic uncertainty, can be imposed on each parameter to penalize it from getting too far from its default value. The following subsections elaborate on the specific choice of measurement observables and model parameters in this work.
4.1 Data points
The observables for the reported differential cross sections across the four TKI measurements—T2K-$0\pi$, T2K-$\pi^{+}$, MINERvA-$0\pi$, and MINERvA-$\pi^{0}$—are detailed in Table 3. To identify the most sensitive variables, various combinations were evaluated for tuning purposes. A total of 26 combinations were assessed, with the superset comprising $\delta\alpha_{\textrm{T}}$, $\delta p_{\textrm{T}}$, $\delta\phi_{\textrm{T}}$, $p_{\textrm{N}}$, and $\delta p_{\textrm{TT}}$ (see Table 6 in Appendix A for all combinations). Each combination was evaluated such that non-selected variables—including proton momentum and angle ($p_{\text{p}}$ and $\theta_{\text{p}}$vA:2018hba , as well as $\delta p_{\textrm{Tx}}$ and $\delta p_{\textrm{Ty}}$ MINERvA:2019ope in MINERvA-$0\pi$—were used solely for validation; the model is expected to accurately predict these variables post-tuning. When constructing the combinations, it was observed that $\delta\phi_{\textrm{T}}$ is strongly dependent on neutrino beam energy Lu:2015tcr , which suppresses its sensitivity to nuclear effects. Additionally, correlations exist between various observables, notably between $\delta p_{\textrm{T}}$ and $p_{\textrm{N}}$. Therefore, determining the optimal combination is not straightforward; this study employs a criterion based on the $\chi^{2}$ with the complete (tuned plus validation) observable set (Fig. 5).